Hi, i'm having trouble with one problem for my homework.
It says: If Cathy walks for 2h and rides her bicycle for 1h, she can travel 36 km. If she walks for 2h and rides her bicycle for 2h, she can travel 56km. How fast can she walk? How fast can she ride her bicycle?
I tried subtracting two equations. The first one 2x+ 1y = 36 and the second one is 2x+2y=56. But when I line them up and subtract, it becomes negative.
Do you know how to do this?

Answer by stanbon(75887)

If Cathy walks for 2h and rides her bicycle for 1h, she can travel 36 km. If she walks for 2h and rides her bicycle for 2h, she can travel 56km. How fast can she walk? How fast can she ride her bicycle?
--------------------
Distance Eq: 2w + r = 36 km
Distance Eq: 2w +2r = 56 km
------
Subtract and solve for "r":
r = 20 km/hr (riding rate)
----
Solve for "w":
2w + 20 = 36
2w = 16
w = 8 km/hr (walking rate)
==============================
Cheers,
Stan H.